OpenStudy (anonymous):

find the polynomial function of degree 4 with zeros 2i and 1-i and a constant term of -24

OpenStudy (anonymous):

do you have any thoughts?

OpenStudy (helder_edwin):

your polynomial has complex roots if \(z=a+ib\) is a zero then \(\bar{z}=a-ib\) also is a zero

OpenStudy (anonymous):

Other than the fact that we only have two zeros, so at least one of them has multiplicity >1 I don't know where to start.

OpenStudy (helder_edwin):

just told u

OpenStudy (helder_edwin):

u HAVE (can deduce) the 4 zeros u need

OpenStudy (anonymous):

so b is 2 and a is 1?

OpenStudy (helder_edwin):

no in one case u have \(z=2i\) and in the other \(z=1-i\)

OpenStudy (anonymous):

so then \[(0+2i)(0-2i)\] and \[(1-i)(1+i)\]

OpenStudy (helder_edwin):

yes your polynomial should be \[ \large f(x)=\alpha(1-i)(1+i)(2i)(-2i) \]

OpenStudy (helder_edwin):

what do you use the \(\alpha\) for?

OpenStudy (anonymous):

thanks! is alpha the constant -24?

OpenStudy (anonymous):

I'm not sure, is alpha -24?

OpenStudy (helder_edwin):

do the multiplication. u'll figure it out

OpenStudy (anonymous):

If I factor out 2 and -2 it seems like 24 should be -6, but I get confused with complex numbers

OpenStudy (anonymous):

I mean alpha should be -6

OpenStudy (helder_edwin):

sorry made a mistake \[ \large f(x)=\alpha(x-1-i)(x-1+i)(x-2i)(x+2i) \]

OpenStudy (helder_edwin):

lets see \[ \large f(x)=\alpha(x^2-2x+(1-i)(1+i))(x^2+4) \] you can go on

OpenStudy (anonymous):

so then alpha should be -8?

OpenStudy (helder_edwin):

give me a second

OpenStudy (helder_edwin):

if u dont take into account the \(\alpha\) you get \[ \large f(x)=\alpha(x^4-2x^3+6x^2-8x+8) \]

OpenStudy (anonymous):

so then its -3